# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ Appendix A Missing Proofs
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})\;\geq\;\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\underline{\chi}_{\mathrm{lo}},$
and by the definition of $\underline{\chi}_{\mathrm{up}}$,
$\underline{\chi}_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}).$
Combining these inequalities with $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$ yields
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\underline{\chi}_{\mathrm{lo}}.$
@@
$\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\quad\text{for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}),$
and, by construction, $\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}$. This is precisely condition (ii).
@@
$\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}\qquad\text{and}\qquad\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\ \text{ for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}).$
@@
$\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}.$
@@
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})\;\geq\;\underline{\chi}_{\mathrm{lo}}$
by Lemma A.8 and the definition of $\underline{\chi}_{\mathrm{lo}}$. Therefore
$\underline{\chi}_{\mathrm{up}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\underline{\chi}_{\mathrm{lo}},$
while the existence of $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ with $\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}$ shows that
$\underline{\chi}_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}.$
Hence $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$, which is condition (i). This completes the proof.
A.5 The proof of Lemma 4.1
@@
$\mathbb{P}(M=m):=\frac{m\,\mathbb{P}(M_{0}=m)}{\mathbb{E}[M_{0}]}\qquad(m\geq 1).$
@@
$=\mathbb{E}_{M,Y_{2:n}}\Bigl[\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\end{subarray}}\Bigl(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr)$
$\qquad\qquad\qquad\quad-\mathrm{e}^{\varepsilon}\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\end{subarray}}\Bigl(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr)\Bigr]$
$=\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr]$
$\quad-\mathrm{e}^{\varepsilon}\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr].$
@@ A.6 The proof of Theorem 5
This proof is an adaptation of the FFT error analysis of Gopi et al. Gopi et al. (2021) to our blanket-divergence setting.
@@
See Section C for a discussion of how to rigorously control these error terms in practice.
@@
Appendix B The Computation of Shuffle Indices
B.1 Generalized Gaussian Mechanisms
Here, we consider the generalized Gaussian mechanism (i.e., Definition 2.6).
Lower Shuffle Index
From the definition, the variance term and the blanket mass in the lower shuffle index are
@@ B.1 Generalized Gaussian Mechanisms
$\sigma_{\mathrm{BG},\beta}^{2}(x_{1},x_{1}^{\prime}):=\gamma_{\beta,\alpha}\left\{\int_{-\infty}^{1/2}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{1}^{\beta}(y)}\,dy\;+\;\int_{1/2}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{0}^{\beta}(y)}\,dy\right\},$
@@
$\gamma_{\beta,\alpha}=\frac{1}{\Gamma(1/\beta)}\int_{(1/(2\alpha))^{\beta}}^{\infty}t^{1/\beta-1}e^{-t}\,dt=\frac{\Gamma\!\bigl(1/\beta,\,(1/(2\alpha))^{\beta}\bigr)}{\Gamma(1/\beta)},$
@@
where $\Gamma(s,x)$ denotes the upper incomplete gamma function. We can see integrands are smooth and rapidly decaying. In particular, for each fixed $\beta>0$ and each pair $(x_{1},x_{1}^{\prime})$, the quantity $\sigma_{\mathrm{BG},\beta}^{2}(x_{1},x_{1}^{\prime})$ (and hence $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$) can be evaluated efficiently by standard one-dimensional numerical quadrature (in the special cases $\beta=2$ (Gaussian mechanism) and $\beta=1$ (Laplace mechanism), the relevant integrals admit closed-form expressions). As a consequence, searching over $(x_{1},x_{1}^{\prime})$ for the lower shuffle index is computationally inexpensive [^11] In all our numerical experiments for generalized Gaussian mechanisms (including the cases $\beta=1$ and $\beta=2$), the pair $(x_{1},x_{1}^{\prime})=(0,1)$ consistently attains (or appears to attain) the infimum of the lower shuffle index $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$. This is also intuitively plausible, since $(0,1)$ maximizes the separation between two inputs in the unit interval. However, a general theoretical proof that $(0,1)$ is the unique global minimizer of $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$ for all admissible parameters $(\alpha,\beta)$ would require a delicate analysis beyond the scope of this work, and we therefore do not claim such a statement as a theorem..
@@
Figure 2(a) illustrates the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ for several values of the shape parameter $\beta$. Even when the underlying local randomizers are calibrated to have the same variance (and hence the same accuracy for mean estimation), they can exhibit different shuffle indices.
@@
Upper Shuffle Index
For the upper shuffle index, we instead take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$ as in Su et al. Su et al. (2025), and only the variance is needed in our analysis. From the definition, the variance is
$\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime}):=\int_{-\infty}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{x}^{\beta}(y)}\,dy\;=\;\alpha\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\,F_{\beta}(a_{1},a_{1}^{\prime}),$ (83)

### v3

@@ Appendix A Missing Proofs
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})\;\geq\;\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\chi_{\mathrm{lo}},$
and by the definition of $\chi_{\mathrm{up}}$,
$\chi_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}).$
Combining these inequalities with $\chi_{\mathrm{up}}=\chi_{\mathrm{lo}}$ yields
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\chi_{\mathrm{lo}}.$
@@
$\mathcal{R}_{x^{\ast}}(y)={\mathcal{R}}(y)\quad\text{for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}),$
and, by construction, $\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}$. This is precisely condition (ii).
@@
$\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}\qquad\text{and}\qquad\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\ \text{ for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}).$
@@
$\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}.$
@@
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}$
by Lemma A.9 and the definition of $\chi_{\mathrm{lo}}$. Therefore
$\chi_{\mathrm{up}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}},$
while the existence of $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ with $\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}$ shows that
$\chi_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}.$
Hence $\chi_{\mathrm{up}}=\chi_{\mathrm{lo}}$, which is condition (i). This completes the proof.
A.4 The proof of Lemma 4.1
@@
$\Pr[M=m]:=\frac{m\,\Pr[M_{0}=m]}{\mathbb{E}[M_{0}]}\qquad(m\geq 1).$
@@
$=\mathbb{E}_{M,Y_{2:n}}\Bigl[\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr]-\mathrm{e}^{\varepsilon}\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr]\Bigr]$
$=\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr]-\mathrm{e}^{\varepsilon}\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr].$
@@ A.5 The proof of Theorem 5
This proof is an adaptation of the FFT error analysis of Gopi et al. [16] to our blanket-divergence setting.
@@
See Section D for a discussion of how to rigorously control these error terms in practice.
@@
Appendix B The Interpretation of Blanket Divergence
Let $\mathcal{R}$ be a local randomizer with blanket distribution $\mathcal{R}_{\mathrm{BG}}$ and blanket mass $\gamma$, so that for every $x\in\mathcal{X}$,
$\mathcal{R}_{x}\;=\;\gamma\,\mathcal{R}_{\mathrm{BG}}\;+\;(1-\gamma)\,\mathcal{R}_{x}^{\mathrm{LO}},\qquad\mathcal{R}_{x}^{\mathrm{LO}}(y):=\frac{\mathcal{R}_{x}(y)-\gamma\mathcal{R}_{\mathrm{BG}}(y)}{1-\gamma}.$
We define an extended mechanism $\widetilde{\mathcal{M}}$ as follows (see Fig. 1). User $1$ always outputs $Y_{1}\sim\mathcal{R}_{x_{1}}$ (or $\mathcal{R}_{x_{1}^{\prime}}$ under $x^{\prime}_{1:n}$). For each $i\in\{2,\dots,n\}$, sample an independent coin $B_{i}\sim\mathrm{Bernoulli}(\gamma)$. If $B_{i}=1$, output $Y_{i}\sim\mathcal{R}_{\mathrm{BG}}$; otherwise output $Y_{i}\sim\mathcal{R}^{\mathrm{LO}}_{x_{i}}$ and reveal it together with its index $i$. Let $T:=\sum_{i=2}^{n}B_{i}$ and let $M:=1+T$ be the number of messages routed to the shuffler. The shuffler is applied only to the $M$ messages $\{Y_{1}\}\cup\{Y_{i}:B_{i}=1\}$, while the leftover messages $\{(i,Y_{i}):B_{i}=0\}$ are released without shuffling. Denote the shuffled part of the output by $S_{1:M}\in\mathcal{Y}^{M}$.
Since the leftover part is identical under $x_{1:n}$ and $x^{\prime}_{1:n}$, the hockey-stick divergence of $\widetilde{\mathcal{M}}$ reduces to that of the shuffled part. Conditioning on $M=m$, the shuffled part consists of one sample from $\mathcal{R}_{x_{1}}$ and m1m-1 i.i.d. samples from $\mathcal{R}_{\mathrm{BG}}$ whose order is uniformly randomized. A direct calculation yields the corresponding density
$p_{x_{1}}^{(m)}(s_{1:m})=\frac{1}{m}\sum_{j=1}^{m}\mathcal{R}_{x_{1}}(s_{j})\prod_{k\neq j}\mathcal{R}_{\mathrm{BG}}(s_{k})=\Bigl(\prod_{k=1}^{m}\mathcal{R}_{\mathrm{BG}}(s_{k})\Bigr)\cdot\frac{1}{m}\sum_{j=1}^{m}\frac{\mathcal{R}_{x_{1}}(s_{j})}{\mathcal{R}_{\mathrm{BG}}(s_{j})}.$
Hence, writing
$l_{\varepsilon}(y):=\frac{\mathcal{R}_{x_{1}}(y)-e^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y)}{\mathcal{R}_{\mathrm{BG}}(y)},$
we have
$p_{x_{1}}^{(m)}(s_{1:m})-e^{\varepsilon}p_{x_{1}^{\prime}}^{(m)}(s_{1:m})=\Bigl(\prod_{k=1}^{m}\mathcal{R}_{\mathrm{BG}}(s_{k})\Bigr)\cdot\frac{1}{m}\sum_{j=1}^{m}l_{\varepsilon}(s_{j}).$
Integrating the positive part, this implies the exact identity
$\mathcal{D}_{e^{\varepsilon}}\!\left(p_{x_{1}}^{(m)}\middle\|p_{x_{1}^{\prime}}^{(m)}\right)=\mathbb{E}_{Y_{1:m}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{BG}}}\left[\frac{1}{m}\left(\sum_{j=1}^{m}l_{\varepsilon}(Y_{j})\right)_{+}\right].$
Finally, we uncondition over $M$. In our construction, $T\sim\mathrm{Bin}(n-1,\gamma)$ and $M=1+T$, so for $m\in\{1,\dots,n\}$,
$\Pr[M=m]=\binom{n-1}{m-1}\gamma^{m-1}(1-\gamma)^{n-m}.$
Therefore,
$\mathcal{D}_{e^{\varepsilon}}\!\left(\widetilde{\mathcal{M}}(x_{1:n})\middle\|\widetilde{\mathcal{M}}(x^{\prime}_{1:n})\right)$ $=\sum_{m=1}^{n}\Pr[M=m]\,\mathbb{E}_{Y_{1:m}\sim\mathcal{R}_{\mathrm{BG}}}\left[\frac{1}{m}\left(\sum_{j=1}^{m}l_{\varepsilon}(Y_{j})\right)_{+}\right]$
$=\sum_{m=1}^{n}\left(\binom{n-1}{m-1}\gamma^{m-1}(1-\gamma)^{n-m}\cdot\frac{1}{m}\right)\mathbb{E}_{Y_{1:m}\sim\mathcal{R}_{\mathrm{BG}}}\left[\left(\sum_{j=1}^{m}l_{\varepsilon}(Y_{j})\right)_{+}\right].$
Using the combinatorial identity $\binom{n-1}{m-1}=\frac{m}{n}\binom{n}{m}$, we obtain
$\binom{n-1}{m-1}\gamma^{m-1}(1-\gamma)^{n-m}\cdot\frac{1}{m}=\frac{1}{n\gamma}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}.$
Since the $m=0$ term contributes 0 (because $(\sum_{j=1}^{0}\cdot)_{+}=0$), we may extend the sum to $m=0$ and recognize the binomial distribution $\mathrm{Bin}(n,\gamma)$:
$\mathcal{D}_{e^{\varepsilon}}\!\left(\widetilde{\mathcal{M}}(x_{1:n})\middle\|\widetilde{\mathcal{M}}(x^{\prime}_{1:n})\right)=\frac{1}{n\gamma}\,\mathbb{E}_{\begin{subarray}{c}M\sim\mathrm{Bin}(n,\gamma)\\ Y_{1:M}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{BG}}\end{subarray}}\left[\left(\sum_{j=1}^{M}l_{\varepsilon}(Y_{j})\right)_{+}\right]=\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}\bigl(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}}\bigr).$
Moreover, the original shuffled mechanism $\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}$ can be obtained from $\widetilde{\mathcal{M}}$ by post-processing (discarding the revealed indices and leftover messages and retaining only the fully shuffled multiset of all $n$ messages). By the data processing inequality for the hockey-stick divergence, this yields
@@ Appendix B The Interpretation of Blanket Divergence
$\mathcal{D}_{e^{\varepsilon}}\!\left(\mathcal{M}(x_{1:n})\middle\|\mathcal{M}(x^{\prime}_{1:n})\right)\leq\mathcal{D}_{e^{\varepsilon}}\!\left(\widetilde{\mathcal{M}}(x_{1:n})\middle\|\widetilde{\mathcal{M}}(x^{\prime}_{1:n})\right)=\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}\bigl(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}}\bigr),$
which recovers the blanket-divergence upper bound.
Appendix C The Computation of Shuffle Indices
C.1 Generalized Gaussian Mechanisms
Here, we consider the $\beta$-Gaussian local randomizer on $\mathcal{X}=[0,1]$:
@@ C.1 Generalized Gaussian Mechanisms
$\mathcal{R}(x)=x+Z,\qquad Z\sim g_{\beta,c},\qquad g_{\beta,c}(z):=\frac{\beta}{2c\,\Gamma(1/\beta)}\exp\!\Bigl(-\bigl|\tfrac{z}{c}\bigr|^{\beta}\Bigr),$
so that $\phi_{x}^{\beta}(y)=g_{\beta,c}(y-x)$.
The blanket density is $\underline{\mathcal{R}}(y)=\inf_{x\in[0,1]}\phi_{x}^{\beta}(y)$. Since $x\mapsto|y-x|$ is maximized over $[0,1]$ at the far endpoint,
$\underline{\mathcal{R}}(y)=\begin{cases}\phi_{1}^{\beta}(y),&y\leq\tfrac{1}{2},\\[2.0pt] \phi_{0}^{\beta}(y),&y\geq\tfrac{1}{2},\end{cases}\qquad\gamma:=\int_{\mathbb{R}}\underline{\mathcal{R}}(y)\,dy=2\int_{1/2}^{\infty}\phi_{0}^{\beta}(y)\,dy.$ (84)
C.1.1 Lower Shuffle Index
From the definition, for fixed $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})$, the variance term and the blanket mass in the shuffle index are
$\sigma_{\mathrm{BG},\beta}^{2}:=\gamma\left\{\int_{-\infty}^{1/2}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{1}^{\beta}(y)}\,dy\;+\;\int_{1/2}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{0}^{\beta}(y)}\,dy\right\},$
@@ C.1.1 Lower Shuffle Index
where $\Gamma(s,x)$ denotes the upper incomplete gamma function. For fixed $\beta>0$, the quantity $\sigma_{\mathrm{BG},\beta}^{2}$ (and hence $\chi_{\mathrm{lo}}$) can be evaluated efficiently by standard one-dimensional numerical quadrature. As a consequence, searching over $(x_{1},x_{1}^{\prime})$ for the lower shuffle index is computationally inexpensive [^11] In all our numerical experiments for generalized Gaussian mechanisms (including the cases $\beta=1$ and $\beta=2$), the pair $(x_{1},x_{1}^{\prime})=(0,1)$ consistently attains (or appears to attain) the infimum of the lower shuffle index $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$. This is also intuitively plausible, since $(0,1)$ maximizes the separation between two inputs in the unit interval. However, a general theoretical proof that $(0,1)$ is the unique global minimizer of $\chi_{\mathrm{lo}}$ for all admissible parameters would require a delicate analysis beyond the scope of this work, and we therefore do not claim such a statement as a theorem..
Here, we adopt the working assumption that the worst-case pair is attained at $(x_{1},x_{1}^{\prime})=(0,1)$, and under this assumption, $\chi_{\mathrm{lo}}$ admits closed-form expressions for the Laplace ($\beta=1$) and Gaussian ($\beta=2$) mechanisms when $\mathrm{Var}(Z)=\sigma_{0}^{2}$. For Laplace, we have
$\chi_{\mathrm{lo}}^{(\beta=1)}=\frac{\sqrt{3}}{2}\cdot\frac{t}{\sqrt{1-3t^{2}+2t^{3}}},\qquad t:=\exp\!\left(-\frac{1}{\sqrt{2}\,\sigma_{0}}\right).$
For Gaussian, we have
$\chi_{\mathrm{lo}}^{(\beta=2)}=\frac{1}{\sqrt{2\left(e^{1/\sigma_{0}^{2}}\,\Phi\!\left(\frac{3}{2\sigma_{0}}\right)-3\,\Phi\!\left(\frac{1}{2\sigma_{0}}\right)+1\right)}},$
These formulas make the limiting behavior explicit: $\chi_{\mathrm{lo}}^{(\beta=1)}(0,1)\sim\sigma_{0}/\sqrt{2}$ and $\chi_{\mathrm{lo}}^{(\beta=2)}(0,1)\sim\sigma_{0}$ as $\sigma_{0}\to\infty$, while both decay exponentially fast as $\sigma_{0}\to 0$.
@@
Fig. 5(a) illustrates the lower shuffle index $\chi_{\mathrm{lo}}$ for several values of the shape parameter $\beta$. Even when the underlying local randomizers are calibrated to have the same variance (and hence the same accuracy for mean estimation), they exhibit different shuffle indices.
@@
C.1.2 Upper Shuffle Index
For the upper shuffle index, we instead take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$, and only the variance is needed in our analysis. From the definition, the variance is
$\sigma_{x,\beta}^{2}:=\int_{-\infty}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{x}^{\beta}(y)}\,dy\;=\;c\frac{\beta}{2c\,\Gamma(1/\beta)}\,F_{\beta}(a_{1},a_{1}^{\prime}),$ (85)